FEL Upgrade Project Weekly Brief
April 16-20, 2001

We are proud to report that the FEL Deputy Program Manager, George "the Flash" Neil completed the Boston Marathon on Monday in 3 hrs and 9 minutes giving him a rank of 61 out of 1600 proto-senior citizens who challenged the event.

F. Dylla attended the Spring 2001 Board of Trustees meeting of SURA and presented an update on the FEL program.  We were pleased to have the Chief of Naval Research, Admiral Jay Cohen address the SURA Board
on Monday night.  The CNR discussed the potential interest of ONR in SURA's new initiative in coastal science and gave mention to the existing ONR-SURA relationship for the development of FELs.  We have had three
discussions with the CNR since early February and during all three occasions he passed on his interest in strengthening an FEL development program involving Jefferson Lab and potential partnerships with Navy
labs and industry.

We spent some of the week making preparations for the Jefferson Lab Open House which will be held this Saturday (April 21). This event is held every two years and typically attracts up to six thousand visitors
from the local area.  We view the open house as a very important component for maintaining our positive community relations.

George Neil gave presentations on the Jlab FEL program this week at MIT, the Air Force Institute of Technology in Dayton and at the Stanford FEL Laboratory.  During the visit to AFIT, George had the opportunity to brief managers in the AFRL in Dayton on our pending UV development program with the Air Force.

WBS 3 (Beam Physics):
A broadband-chromatic model was developed and applied to  optimize the energy-recovery endloop betatron matching.  The  resulting solution provides space for Thomson scattering x-ray  generation apparatus.

Field quality issues in the pi-bends and injection line dipoles were under discussion.

WBS 4 (Injector):
Design work continued on the region between the gun and the lightbox, we are working to maximize the aperture for cathode coverage and  electron beam transport. Design changes for the ball cathode were completed. Ordered bellows for the stalk and shield modifications

WBS 5 (SRF):
All the helium vessels have been delivered and passed their QA tests.  We have completed the electron beam welding for two fundamental power couplers (FPC) bodies and qualified the procedure. A kick-off meeting for the thermal shield vendor (ADC) has been scheduled for April 25.

Cavity vertical test assembly (VTA) test began last week for the 7-cell cavity structures. Two test have been completed after the standard buffered chemical polish (BCP) treatment.  These cavities show onset of
field emission at 11 MV/m. They will be re-treated and re-tested until higher gradients are achieved.

WBS 6 (RF):
Continued maintenance work on Quarter power supply.
Pursuing specifications for transformer and inductors for Quarter.

WBS 8 (I&C):
Rip out meeting are being held and there is good progress on progress on planning the transition from
1(+)kWatt to >10kWatts.

Cable tray has already been ordered. LCW and vacuum layouts/requirements are also progressing well.

The Laser Safety System (LSS) master has been updated for operating multiple hutches in any or all of the user labs. There is now available status for operation of up to 5 hutches in the six labs. An "FEL beam-block" shutter has been developed by the optics group that will allow each hutch to be manually bypassed for rapid access and sample changes without dropping the labs LSS.  This beam-block and the user lab shutter provide the two independent methods of preventing severe biological damage from the FEL beam to the user. The Lab 2 LSS is installed and complete. Final certifications for all of the LSS components require the FEL to be operating so this will be done just prior to the summer run. (Note: preliminary certs have been done as a first check and as a training exercise)

Parts for a prototype BPM using Motorola DSP56F807 and serial CAN bus including filters and down-convertors have been ordered. The software development kit is in for this DSP. The BPMs are a couple weeks late from Modern Machine but do not yet threaten the schedule. 20 feet of table space has been setup in the back of the FEL for the electrical assembly of the beam viewers prior to passing them off for "girder" installation. During the last go-round this was done in the tunnel - we're learning.

The new design for the OBPM PCB using the Molectron detectors is complete and assembled, once the open house is over this will be tested. Mike has been working on chassis and rack drawings to add to the web site.  He also is detailing the machining for the boxes for additional cameras that will be built.

Ion Pump Power Supply documentation continues.  Drawings are back for the Charge/Dump Current Monitor Chassis.  Jim Coleman is assisting us to get a system drawing for the Machine Protection System together.  James, Mike, Dan and Deborah were trained on the correct procedures to perform a sweep of the machine tunnel so it can be taken to "beam permit".

Drawings currently at EECAD:  Charge Totalizer Board; Picomotor Controller Chassis Wiring Diagram; LSS Master Chassis Wiring Diagram and LSS User Lab Interface Box Wiring Diagram.

Everyone has been straightening up the labs and sprucing up the place in preparation for the open house this weekend.

WBS 9 (Transport):
Injector Dipoles (DU/DV)
o  DULY Research is still trying to get their magnetic model of the small injector Dipole (GV) to achieve the
    required tolerances of 1 part in 10,000 for both field flatness and field integral.  The problem is that they are
    mutually exclusive for most of the tricks they have tried thus far.  At weeks end they have achieved field
    flatness using a trim coil without altering integral.  Altering the magnet face into a curved surface altered field
    integral strongly.  They are working at combining the two tricks.
o They have completed drafts of the details of the GU that are able to be checked by us.
Optical Chicane Dipole (DW)
o The entire back checked package was rechecked and given back to the designer to make some small
Arc Dipoles (GY, GX, GQ)
o  AES continues progress in designing the 180 degree dipole (GY). They added fillers under the coils for missing
    pole tip material and added a preliminary hole for the periscopes.  They will hold off on any further changes
    pending resolution of the magnetic model's final solutions.
o  The magnetic analysis of the GY with path length correctors continued.  At weeks end Tom Schultheiss is
    calculating parallel offset if a beam is brought into the magnet perpendicular but off set and comes out in a
    similar condition and the bulk field is the required field.  The face of the magnet needs extension through these
    iterations and it looks like the offset will be about 7 mm.
o At the same time he is modeling the two sets of periscopes holes through the yoke plates, near the coils at
    around 45 degrees in the GY.
o They will begin design of the GX next week.
QG (3 inch quad)
The coil manufacturing package is in Procurement.
QX (Trim Quad)
o We continued to debug our magnetic model.
Corrector Dipoles (DB & DJ)
o The low bid was obtained from a new vendor.  We are going to their factory on Monday to assess if it is a
    qualified bid.
o We set up a contract with a consulting firm (Innovation Services) to evaluate our epoxy insulation scheme and
    the manufacturing specifications for the dipoles.
o The Dipole Steel Procurement Package was signed and went to Procurement.
o We started making the tradeoffs of machined in features vs adjustment mechanisms in our laying out and design
    the girders and stands.
o We continued firming up the parameters associated with dipole chamber height and the configurations of the arc
o We are sending out inquires to vacuum vendors to determine if they can make our style of chamber.  The Jlab
    Shop, which did much of the chamber work last time, may be overloaded at our required time of fabrication.
o We took our first look at piping the LCW in the upgrade.

WBS 11 (Optics):
The optical cavity assembly design progresses, it now has a stiffer support frame and accommodates an exterior motor for translation.  The check prints for the new ultraviewer will be produced today.

Some effort was spent preparing several of the User Labs for the Lab Open House, and helping users store their equipment. The manually-actuated turning mirror (initially for the User Lab 3) was submitted to our machine shop for fabrication.  I & C was supported in their efforts to test out the LSS.  The LSOP for User Lab 2 has been signed off within the FEL Dept., and is now with Safety Dept. officials.  I gave a talk at the Dartmouth College Physics and Astronomy Dept. Colloquium.

No operations were scheduled this week.   A call for proposals for the planned summer user run was distributed by e-mail to the user list.